CIT 3322 CSC 3322 Computer Architecture

This is designed to comprehensively clarify the number systems. This assignment weights 10 points.

Note:

A cover page with "CIT 3322 CSC 3322, Spring 2019, Assignment #1", your

NAME

, and

your

E-MAIL address

on it.

My preference is typed report. Writing and hand-draw are OK but it should be legible.

Staple if there are more than one sheet.

You can discuss with your classmate, but don't copy.

1. Convert the following decimal numbers to (regular) binary numbers:

a. 99

b. 48.25

2. Convert the following (regular) binary numbers to the decimal numbers:

a. 101010

2

b. 110.10101

2

3. Convert the following negative decimal numbers to binary numbers using three

different kinds of binary representations (sign-magnitude, 1’s complement, and 2’s

complement), respectively:

a. -28

b. -52

4. Convert the following 2’s complement binary numbers to decimal numbers:

a. 010101

b. 101010

5. Covert the following decimal number to BCD code:

a. 49

b. 136

6. Convert the following binary code to gray code:

a. 10101

b. 1010111

7. Convert the following gray code to binary code:

a. 11111

b. 1111100
Answer #1

1 a)1100011

b)110000.01

2. a) 101010 = (1 × 2⁵) + (0 × 2⁴) + (1 × 2³) + (0 × 2²) + (1 × 2¹) + (0 × 2⁰) = 42

b)110.10101 = (1 × 2²) + (1 × 2¹) + (0 × 2⁰) + (1 × 2⁻¹) + (0 × 2⁻²) + (1 × 2⁻³) + (0 × 2⁻⁴) + (1 × 2⁻⁵) = 6.65625

3 a) -28=1111111111100100 //binary 2's complement

b) -52=1111111111001100 //binary 2's complement

4 a) 21

b) -22

5) a) 49 = 0100 1001 //In the BCD numbering system, a decimal number is separated into four bits for each decimal digit within the number.

b) 136 = 0001 0011 0110

6) a) 10101=11111

b) 1010111 = 1111100

7) a)10101

b)1010111